please help me, this is the last problem i couldnt solve for my assignment.

A FACTORY MEASURING 80m BY 60m IS TO BE SURROUNDED BY A SIDEWALK UNIFORM IN WIDTH. IH THE AREA OF THE FACTORY IS DOUBLED BY EXTENDING ALL SIDES, WHAT IS THE WIDTH OF THE SIDEWALK (ROUND TO THE NEAREST TENTH)?
thank you

Answer by checkley71(8403)

THE ORIGINAL FACTORY HAS A PERIMETER=2*80+2*60=160+120=280 CM.
THE WIDTH OF THE SIDEWALK IS NOT KNOWN/GIVEN FOR THE ORIGINAL FACTORY.
THE EXPANDED MEASUREMENTS OF THE FACTORY ARE 2*160+2*120=320+240=560 CM.
THEREFORE THERE IS NO WAY OF CALCULATING THE WIDTH, ONLY THE EXPANDED LENGTH OF THE SIDEWALK.